# Interpreting the power spectral density of a fluctuating colloidal current

### v1

@@ III.4 Developing a distributed transit time model for the PSD
While the overall form of our experimental current spectra is similar to the transit time model, the two curves do not agree quantitatively using the mean transit time as input. Here, an obvious discrepancy is that the model assumes a single transit time, whereas in the experiment there is a distribution of particle transit times due to the variation of fluid flow speed across the channel width. This suggests that to establish a fully quantitative model, Eq. 7 must be extended to account for the particle transit time distribution.
The first ingredient determining the shape of this distribution is the velocity profile of particles within the channel. This can be calculated directly from the experimental trajectories Knowles et al. (2022) and a typical example is shown in Fig. 4(a). Particle velocities inside the channel are dominated by advection and so reflect the underlying fluid flowGuazzelli and Morris (2011), with corrections for finite particle size and hydrodynamic interactions with the confining walls Chen and Ye (2000); Pasol et al. (2011). As such, particle velocities are lowest at the channel walls due to friction and rise towards the channel centre, similar to a Poiseuille flow profile.
Figure 4: Distributions of particle position and speed inside the channel. (a) The particle velocity along the channel as a function of distance from the channel centre. Data points show the mean speed with $\pm$ one standard deviation as error bars. (b) The normalised probability distribution of particle positions across the channel. (c) An illustration of the particle capture process under pressure driven flow at high Peclet number. (d) The distribution of particle speeds within the channel, $P(v)$.
Figure 5: A comparison between measured (points) and modelled (lines) power spectral densities of colloidal current fluctuations. Rows have constant flow speed, indicated in the leftmost panel, and columns have the same ROI length, indicated at the top of the column. Insets in rightmost panels show the velocity distribution for the three different flow speeds.
The probability distribution for the speed of a randomly selected particle does not exhibit this same functional form, however, as particle streamlines are not uniformly distributed across the channel width. As such, the probability distribution for particle position within the channel, $P(y)$, must also be accounted for. $P(y)$ corresponding to the velocity distribution in (a) is plotted in Fig. 4(b) and shows an excess of particles at the channel walls. This is an excluded volume effect. If a particle enters the channel on a fluid streamline that lies within one particle radius of the channel wall, the particle must be shunted closer to the channel centre, as shown schematically in Fig. 4(c). This creates an enhanced probability of finding particles in a small region close to the walls.
Previous work observed a similar effect, where mean velocities of tracer particles in microchannels were lowered by a bunching of particles at the channel wall Liot et al. (2018). Here, the authors introduced a dimensionless entrance Péclet number, comparing the time to be advected into the channel and the time to diffuse across half its width. Values of $Pe>\sim 10^{3}$ are associated with fast capture at the entrance to the channel, meaning particles have insufficient time to redistribute themselves diffusively across channel streamlines. Calculating this quantity in our experiments gives values $Pe\sim 2\times 10^{3}$ and higher, consistent with this previously predicted regime Liot et al. (2018).
The probability density function for the speed of a randomly chosen particle passing through the channel, $P(v)$, follows from $v(y)$ and $P(y)$, and is shown in Fig. 4(d). Here, the probability density $P(v)$ was obtained directly as a normalised histogram of individual particle velocities within the channel, but can be qualitatively interpreted from the two distributions in panels (a) and (b). In particular, the main peak in $P(v)$ is at the speed realised in the centre of the channel and the tail at lower velocities comes from the excess of particles at the edge. In the following, we choose the probability distribution of particle velocities $P(v)$ as the most useful experimental measure of the distribution of particle transit times; compared to a distribution of transit times, it is independent of ROI length and has a finer resolution than measuring channel occupancy times, which are coarsened by the finite frame rate.
@@
To extend Eq. 7, we now proceed by treating the total power spectrum as the sum over power spectra from many independent sources, where each source has a different characteristic transit time arising from the velocity distribution. Firstly, we define $P(\tau)$ to be the probability density function that a particle crosses the ROI in time $\tau$. The contribution to the current comprised of particles transiting in time $\tau$ is then:
$I_{\tau}=I_{0}P(\tau)d\tau,$ (8)
@@
This is a complete model for the PSD of shot noise fluctuations in a current with a distribution of transit times. If we transform this into the particle speed domain, which is experimentally more convenient, we have $\tau=L/v$ and $P(\tau)d\tau=P(v)dv$
@@
The distributed transit time model retains many features of the single transit time model. Firstly, there is a corner frequency $f_{c}=1/(2\tau)$, now at a characteristic average transit time, separating $f^{0}$ and $f^{-2}$ noise regimes. The $f^{0}$ scaling arises due to the independent random arrival of particles, while the $f^{-2}$ scaling arises from the linear scaling in the autocorrelation function, a consequence of the uniform velocity of particles as they transit across the ROI. The key difference is that oscillations are smoothed out in the distributed transit time model. This smoothing is sensitive to the distribution of particle transit times. For example, a wider distribution of transit times (or a narrower distribution of particle speeds) creates more smoothing out of these oscillations.
Finally, in Fig. 5 we show a comparison between the model spectrum of Eq. 11 (solid line) and the experimental data (points) for three different flow speeds and three different values of $L$. In all cases, the modelled spectrum was evaluated from Eq. 11 as a sum over the experimental distributions for $P(v)$ (shown as insets) with the mean current $I_{0}$ calculated directly from the measured particle currents. Excellent agreement is seen in all cases across four decades of frequency, especially in the size and location of the oscillations at around 1 Hz for larger ROIs. The corner frequencies are also a good match to $f_{c}=1/(2\tau)=\bar{v}/(2L)$, using the mean particle speed as input. At low frequencies, the theoretical expression appears to slightly overestimate the height of the plateau, and the data rises above the model slightly at the highest frequencies due to aliasing of the spectrum Kirchner (2005). We emphasise, however, that in comparing Eq. 11 to the experimental data here, there are no free fitting parameters: the only inputs are the defined ROI length, $L$, along with the average current, $I_{0}$, and the speed distribution, $P(v)$ which are measured directly from the particle trajectories.
@@ IV Discussion and Conclusions
In this paper, we have explored the fluctuation behaviour of low-density currents of colloidal particles driven through confining channels in a microfluidic device. Inspired by experiments on nanopore transport, we quantity fluctuations by calculation of the power spectral density (PSD) of the current with the goal to link features of the PSD to underlying physical mechanism. Despite the relative simplicity of the experiment, the power spectral density of the current fluctuations through the channel shows rich behaviour, with multiple characteristic features. For this well-defined and observable experimental system, however, we can unambiguously link these features to specific aspects of the colloidal dynamics as follows.
  • The PSD shows two characteristic scaling regimes separated by a corner frequency, $f_{c}$, which reflects the time it takes a particle to cross the ROI. As such, the high frequency scaling regime relates to times that are short compared to this transit time and reflects transport of the particle through the channel itself. In contrast, the low frequency regime relates to a fluctuation mechanism operating over much longer time scales. An obvious candidate for the low frequency behaviour is fluctuations in the entrance of particles into the channel.
  • At low frequencies, we find a scaling as $f^{0}$, implying that arrival times of particles into the channel are uncorrelated. We have previously shown that particle transport into the channel depends upon the time for drift-diffusion across the reservoir to a capture regionKnowles et al. (2022). For the semi-dilute packing fractions studied here, the mean distance between particles is relatively large and so interparticle interactions do not significantly influence dynamics. As such, particles do not impede each other and enter the channel independently, consistent with our observation of low frequency white noise.
  • The amplitude of the low frequency regime is approximately twice the average current through the ROI. This is the low frequency limit predicted for shot noise.
  • At high frequencies, we see a decay as $1/f^{2}$ as the finite time for particles to pass through the channel leads to time-dependent correlations in the particle current. Observation of a scaling of $1/f^{2}$ reflects the directed, advective transport of particles. This is distinct from diffusive transport through channels Marbach (2021), which been observed to lead to an alternative scaling of $1/f^{3/2}$.
  • The high frequency regime exhibits characteristic oscillations that arise from the finite transit time of particles through the channel. The extent to which these are observed depends on the distribution of transit times, a consequence of the distribution of particle velocities within the channel. Inside our microfluidic devices, the distribution of transit times has a complex functional form that can be rationalised by consideration of the nonuniform velocity profile across the channel width and nonuniform distribution of particle positions. In general, however, oscillations in the PSD are more or less pronounced depending on the width of the distribution, with a narrower range of transit times giving larger amplitude oscillations.
@@
This combination of PSD features is reminiscent of models for shot noise with a finite transit time developed for electronic noise. Despite the obvious physical differences between the two systems, we have shown that by modifying such expressions to allow for a distribution of finite transit times we can quantitatively model the experimental power spectral density of our colloidal currents. This highlights firstly how widespread these particular noise creating mechanisms are across physical systems and secondly the potential for translating models from the well-studied area of electronic noise to understand fluctuations in soft systems. Moreover, comparison to these shot noise models demonstrates that while the characteristic scalings of the PSD can be explained by rather general mechanisms, more specific details of the system, for example, the velocity distribution arising from our specific microfluidic geometry, are required to achieve a full quantitative understanding.
Our experimental model is heavily simplified in comparison to synthetic and biological nanoscale pores, but its key features are still relevant. Molecular currents are also comprised of randomly arriving current-carrying particles and as such similar signatures in the PSD could in principle be observed for molecular transport. Moreover, while the range of transit times in the colloidal model was due to the particle velocity distributions as we worked in an advection-dominated regime, distributions of transit times due to stochastic variation in diffusive transport could be explored in a similar manner.
More broadly, having fully and unambiguously established the behaviour of the power spectra for the simple case of a non-interacting current of hard spheres, this experimental platform offers an opportunity to investigate the impact of more complex phenomena on the PSD. For example, in highly crowded molecular systems capture of particles may no longer be independent, which would require an extension of the analytical model to account for interactions. Moreover, many open questions remain about the origin of low frequency $f^{-1}$ scaling of the fluctuations in nanopores. Recent work has suggested that this scaling arises as a consequence of particle dynamics in the reservoir rather than in the channels themselves Robin et al. (2023a); Fragasso, Pud, and Dekker (2019) and resolving such behaviour in our colloidal model will be the object of future work.

### v2

@@ III.4 Developing a distributed transit time model for the PSD
An obvious source of the discrepancy between experiment and model is that the model assumes a single transit time, whereas in the experiment there is a distribution of particle transit times due to the variation of fluid flow speeds across the channel. We quantify this from measured trajectories via the probability distribution of particle velocities $P(v)$, where the distribution is calculated over the ensemble of each particle’s mean velocity. Compared to a distribution of transit times, $P(v)$ is independent of ROI length and has a finer resolution than measuring channel occupancy times, which are coarsened by the finite frame rate. This makes the velocity distribution a more convenient experimental parameter.
@@
To extend Eq. 5, we proceed by treating the total power spectrum as the sum over power spectra from many independent sources, where each source has a different characteristic transit time arising from the velocity distribution. Firstly, we define $P(\tau)$ to be the probability density function that a particle crosses the ROI in time $\tau$. The contribution to the current comprised of particles transiting in time $\tau$ is then:
$I_{\tau}=I_{0}P(\tau)d\tau,$
@@
Finally, transforming this into the particle speed domain, which is experimentally more convenient, we have $\tau=L/v$ and $P(\tau)d\tau=P(v)dv$, such that:
@@
Figure 4: Building the experimental probability distribution of the velocity from underlying physical factors. In all cases data is shown for two mean particle velocities, as indicated in panel (b). (a) The particle velocity along the channel as a function of distance from the channel centre. Data points show the mean speed with $\pm$ one standard deviation as error bars. Solid lines show the modelled profile for Poiseuille flow in a rectangular channel. (b) The corresponding experimental probability distributions of particle velocity, $P(v)$ (solid lines, left axis), compared to the model for $P(v)$ obtained from $v(y)$ and a flat $P(y)$ (dashed line, right axis). (c) The normalised probability distributions of particle position across the channel, corresponding to the velocity profiles in (a). (d) A comparison of data and the model for $P(v)$ determined from $v(y)$ and experimental $P(y)$. (e) The experimental probability distribution of the velocity at the channel centre, $P(v|\delta y)$ (points), compared to the predicted distribution from modelling fluctuations in particle height. Mean particle speeds for this range of positions are indicated in the legend. Inset shows the predicted distribution of particle positions away from the base of the sample cell. (f) A comparison of experimental probability distributions of the velocity with the modelled $P(v)$ determined from models for $v(y)$ and $P(v|\delta y)$.
The experimental probability distribution of the velocity, $P(v)$, would not be accessible in a nanopore experiment. This quantity can, however, be directly measured from particle trajectories in our system, allowing us to assess if including $P(v)$ in Eq. 6 is sufficient to model our experimental spectra. To this end, in Fig. 3(b) and (c) we show an equivalent comparison between the experimental data and model spectrum of Eq. 6, using $P(v)$ (insets) as input. Excellent agreement is seen across four decades of frequency, especially in the size and location of the oscillations at high frequency, for two different mean particle velocities. Similarly good agreement is seen for a range of box sizes as shown in Supplementary Information, Fig. LABEL:fig:Pvy_allspeeds. Importantly, in comparing Eq. 6 to the experimental data there are no free fitting parameters: the only inputs are the defined ROI length, $L$, along with the average current, $I_{0}$, and the probability distribution of the speed, $P(v)$ which are measured directly from the particle trajectories. This indicates that in this experimental scenario, all relevant fluctuation mechanisms are captured by $I_{0}$ and $P(v)$. While the value of $I_{0}$ can be simply rationalised in terms of particle flux through the system, it is clear from the insets of Fig. 3 that the probability distribution of the velocity has a complex shape. We now consider the factors that govern the experimental $P(v)$ and seek an analytic model for this quantity.
III.5 Modelling $P(v)$ for a colloidal current
To first order, $P(v)$ arises from two underlying features: the velocity profile in the channel, $v(y)$, and the probability of particles transiting along a particular streamline with this velocity, $P(y)$. These quantities can be combined to find $P(v)$ as:
$P(v)=P(y)\frac{dy}{dv}.$ (7)
The velocity profile of particles within the channel, $v(y)$, can be calculated directly from the experimental trajectories Knowles2022 and typical examples at high and low mean velocities are shown in Fig. 4(a). Particle velocities inside the channel are dominated by advection and so reflect the underlying fluid flowGuazzelli2011 , with differences due to the finite size of particles and their hydrodynamic interactions with the confining walls Chen2000 ; Pasol2011 . As such, particle velocities are lowest at the channel walls due to friction and rise towards the channel centre as in a pipe flow profile. Solid black lines in Fig. 4(a) show the flow profile obtained via a numerical solution of Poiseuille flow for a rectangular channel with a first order correction for the finite particle diameter. To obtain this, we use the known height and width of the experimental channels, assume that particles move in a plane $\sim\sigma/2$ from the base of the cell and rescale the final curve by the measured mean particle velocity (see Supplementary Information Section LABEL:sec:velocity_profile). Overall, the shape of the velocity profile across the channel is in good agreement with the predicted flow profile, with same small deviations in both profiles likely coming from some variation in the channel height.
Assuming that particles are evenly distributed across the channel width, i.e. $P(y)$ is constant aside from the excluded volume at the walls, Eq. 7 can be used with the numerical solution for $v(y)$ to obtain a first approximation to $P(v)$. This is shown as a black dashed line in Fig. 4(b) where it is compared to the experimental distributions. Despite the good agreement of the velocity profiles, poor agreement is seen between the experimental velocity distributions and this model. In particular, sampling the velocity profile, $v(y)$ gives a distribution of particle speeds that is heavily skewed to faster particles. This can be seen from the shape of the profile in which the top $50~{}\%$ of possible speeds accounts for $\sim 73~{}\%$ of the channel width, such that most of the distribution’s weight is found at a relatively narrow set of speeds.
Figure 5: A comparison between experimental (points) and modelled (lines) power spectral densities. Modelled spectra in the left, middle and right columns are derived from different approximations to the experimental $P(v)$, shown as black dashed lines in Fig. 4(b), (d), and (f) respectively. Rows have constant flow speed, indicated in the leftmost panel, and all data is shown for $L=64\mu$m.
From Eq. 7, a second contributing factor to $P(v)$ is the probability distribution of particle positions in the channel, $P(y)$. Distributions corresponding to the velocity profiles in Fig. 4(a) are plotted in panel (c). Note that here in calculating the distribution, we consider the distribution of positions that have first been averaged over each particle trajectory, such that $P(y)$ provides a measure of the probability of a given particle following a particular streamline. Over the central region of the channel, $P(y)$ is very similar for the two flow speeds, but $P(y)$ shows an excess of particles at the channel walls for higher mean particle velocities.
As the time for a particle to diffuse across the channel width is much longer than the particle transit time, particles have insufficient time to redistribute themselves diffusively across channel streamlines. As such, it is likely the shape of these distributions is dominated by the way in which streamlines in the reservoir are squeezed upon entering the confining channel. In particular, if a particle enters the channel on a fluid streamline that lies within one particle radius of the channel wall, the particle must be shunted closer to the channel centre, creating an enhanced probability of finding particles in a small region close to the walls for high velocities. Previous work observed a similar effect, where mean velocities of tracer particles in microchannels were lowered by a bunching of particles at the channel wall Liot2018 . Here, the authors introduced a dimensionless entrance Péclet number, comparing the time to be advected into the channel and the time to diffuse across half its width. Values of $Pe\gtrsim 10^{3}$ are associated with fast capture at the entrance to the channel, and the values $Pe\sim 2\times 10^{3}$ and higher in our experiments are consistent with this previously predicted regime Liot2018 .
To determine whether the shape of $P(y)$ is a significant factor in determining $P(v)$, in Fig. 4(d) we show the prediction for $P(v)$ from Eq. 7 derived from the experimental $P(y)$ and numerical solution for $v(y)$. It is clear from this comparison that including a more accurate $P(y)$ only slightly improves agreement with the experimental data. This indicates that an additional source of fluctuations in our data is important.
While a Poiseuille flow model for $v(y)$ provides a good prediction for the mean particle velocity, it is clear from the error bars in Fig. 4(a) that the variation in velocity for particles at a particular distance from the channel walls is relatively large considering the high Peclet numbers associated with particle dynamics in the channel ($Pe>27$)Knowles2022 . To understand the origin of this, in Fig. 4(d) we show the probability distribution of the velocity for the subset of particles transiting the channel close to the central axis (denoted as a grey shaded region in (a)). For this range of positions in $y$ the mean velocity varies by only $0.5\%$, however, $P(v|\delta y)$ illustrates that particles clearly exhibit a much broader distribution of velocities. The variation in velocity here shows no dependence on time or distance travelled along the channel and is too large to be accounted for by tracking error or polydispersity. Additionally, while the channel is often occupied by more than one particle, particle velocity shows no dependence on channel occupancy, in line with previous findings for pressure-driven flowMisiunas2019 .
Instead, this variation in velocity appears to arise from very small fluctuations in the distance of a particle from the base of the sample cell. Such fluctuations in $z$ on the order of tens of nanometers are too small to be resolved from the microscopy images, but can lead to significant velocity fluctuations due to the steep variation of the velocity profile close to a wall. To assess the magnitude of such fluctuations, we follow Alexander and Prieve Alexander1987 and estimate the probability distribution of particle positions in $z$ by balancing gravitational and electrostatic potentials (see Supplementary Information Section. LABEL:sec:vsmear). This distribution is shown inset in Fig. 4(e) and to a first approximation is independent of driving force. $P(z)$ can then be combined with the modelled velocity profile away from the wall to obtain a prediction for $P(v|\delta y)$. This is shown as a black dashed line in Fig. 4(e) and is in excellent agreement with the experimental distribution for $\langle v\rangle=36.1\mu$ms-1. Our model for $P(v|\delta y)$ does not predict a dependence on the overall magnitude of the particle velocity when rescaled as in Fig. 4(e). As such, the fact that we observe a slightly broader distribution for slower mean particle velocity could indicate speed-dependent effects like a small hydrodynamic lift force, or perhaps the onset of diffusive effects for lower Peclet number. We note, however, that the width of this distribution is, in general, relatively insensitive to mean speed as shown in Supplementary Information Fig. LABEL:fig:Pvy_allspeeds.
Finally, in Fig. 4(f) we compare the experimental data to a model for $P(v)$ that samples the numerical solution for the velocity profile, $v(y,z)$, across the whole channel cross-section according to our model for $P(z)$. For simplicity, we here neglect any variation of $P(v|\delta y)$ with mean particle velocity of distance from the channel wall (more details are provided in Supplementary Information, Section S2). Despite this approximation, good agreement is now found between the experimental data and model, indicating that we are now capturing the key physical phenomena governing $P(v)$.
III.6 Sensitivity of the power spectral density to the underlying probability distribution of the velocity
Having established a minimal model to describe the behaviour of $P(v)$, we now explore how important the shape of this underlying distribution is to the features of the power spectral density. Fig. 5 shows a comparison between the experimental spectra and the PSD obtained from Eq. 6 using the three models for $P(v)$ from Fig. 4(b), (d) and (f) (marked as I, II and III, respectively). It is clear that just including the velocity distribution for Poiseuille flow significantly improves the agreement between experimental data and modelled spectra compared to a single transit time model. The high frequency oscillations are observably slightly too sharp for this case, however, particularly when considering behaviour across a range of box sizes and driving forces. (see Supplementary Information Section LABEL:sec:allspectra). In line with the behaviour of $P(v)$, including the distribution of particle positions (model II) does not significantly improve the modelled PSD . In contrast, considering the effect of fluctuations in particle height on the distribution of velocities produces model spectra (III) that are essentially indistinguishable from the PSD obtained from the full experimental $P(v)$, as shown in Fig. 3.
@@ IV Discussion and Conclusions
In this paper, we have explored the fluctuation behaviour of low-density currents of colloidal particles driven through confining channels in a microfluidic device. Inspired by experiments on nanopore transport, we quantify fluctuations by calculation of the power spectral density (PSD) of the current with the goal of linking features of the PSD to underlying physical mechanisms. Despite the relative simplicity of the experiment, the power spectral density of the current fluctuations through the channel shows rich behaviour, with two distinct scalings for high and low frequency regimes. These link to fluctuations in particle transport through the channel and fluctuations in particle entrance to the channel, respectively. Our experimental spectra are reminiscent of those for shot noise with a finite transit time, which has been modelled extensively in the context of electronic noise. Despite the obvious physical differences between the two systems, we have shown that we can use such expressions to quantitatively model the experimental power spectral density of our colloidal currents if they are modified to include a distribution of transit times. This highlights, firstly, how widespread these particular noise-creating mechanisms are across physical systems and, secondly, the potential for translating models from the well-studied area of electronic noise to understand fluctuations in soft systems.
While comparison to shot noise models demonstrates that characteristic scalings of the PSD can be explained by rather general mechanisms, such mechanisms can only take our understanding so far. Fully resolving trajectories in this platform has allowed us to demonstrate how subtle details of the spectra quantitatively depend on specific physical properties of the system, in this case the distribution of particle velocities. Again, even for this relatively simple experimental scenario, particle velocities might depend on many potential factors. The high level of control inherent to our system enabled us to pinpoint the dominant factors contributing to the distribution, namely flow profiles of the suspending fluid in the confined geometry and fluctuations of particles with respect to these. Importantly, it also makes it possible to exclude alternative explanations, e.g. linked to interparticle interactions or spatial variations in channel shape. This powerful trajectory-based approach is inaccessible for nanopore experiments for which only an integrated current can be measured and where many physical details are not precisely known.
Although this experimental model is clearly heavily simplified in comparison to synthetic and biological nanoscale pores, our colloidal analogue allows us to understand mechanisms relevant to the latter.Random, independent particle arrivals to the pore, as well as a distribution of through-pore transit times, are both factors which could influence molecular transport. Our work demonstrates quantitatively how these mechanisms manifest in the PSD for transport phenomena with specific physical properties. In linking to molecular studies, our experimental geometry enables observation of behaviour in both the reservoir and the channel. This is crucial for understanding nanopore spectra as, for these finite size channels, the low frequency regime of the PSD measured inside the channel reflects particle entrance statistics; a process that is governed by behaviour outside the channel itself.Robin2023 ; Fragasso2019 Such phenomena can thus not be explored in studies that only consider the confined channel transport dynamics in isolation.
More broadly, having fully and unambiguously established the behaviour of the power spectra for the simple case of an effectively non-interacting current of hard spheres, this experimental platform provides a valuable opportunity to investigate the impact of more complex phenomena on the PSD. Many open questions remain about the origin of low frequency $1/f^{\alpha}$ scalings ($20<\alpha<$ 2) in the power spectra for nanopores. Interestingly, despite the ubiquity of $1/f^{\alpha}$ scalings for nanopores, our colloidal system as realised shows no evidence of $1/f^{\alpha}$ noise at low frequency. Moreover, having fully established the mechanisms underlying the low frequency fluctuation behaviour in our experiment, there appears to be no phenomena that would induce $1/f^{\alpha}$ noise for this colloidal system even over longer time scales, assuming a steady state is maintained. The absence of $1/f^{\alpha}$ noise means that our current system is an ideal starting point from which to explore what additional physical phenomena are required to generate specific low frequency scalings in the PSD. This question has been extensively studied in theory and simulation, but an unambiguous experimental elucidation of this problem is lacking. Investigating these behaviours will thus be the subject of future work.
@@
Supplementary Information
The supplementary information contains details on modelling the velocity profile, $v(y)$ and the velocity fluctuations arising from the distribution of particle heights. Also, additional figures illustrating power spectra at a range of speeds and ROI sizes and the dependence of the velocity distribution $P(v|\delta y)$ on particle position within the channel and driving speed.
Author Contributions
These authors contributed equally to this work.